OpenStudy (anonymous):

the sum of the three angles of any triangle is 180 degree. In a certain triangle, the second angle is 24 degree less than the sum of the other two angles, and the sum of the first and the second angles is 11/7 of the third angle. find the measure of each angle.

OpenStudy (anonymous):

Well let: a=first angle b=second angle c=third angle from your post we can set up a system of equations: (1) a+b+c = 180 (2) b = a+c -24 (3) a+b=11c/7 now we have to solve the system. substitution is the easiest way. substitute (2) into (1) a+[a+c-24]+b+c=180 2a+2c = 180 + 24 2(a+c) = 204 a+c = 102 Now that you know this, plug this into (1) a +b+c=180 b +(a+c) = 180 b + [102]=180 b = 180-102 b = 78 Now put this in equation (3) a+b = 11/c/7 a+78 = 11c/7 a = 11c/7 -78 and throw it in (1) a+b+c =180 11c/7-78 + 78 +c = 180 11c/7+7c/7 = 180 18c/7 = 180 c = 180*7/18 c = 70 Finally now that you know two of the angles: a+b+c=180 a+78+70=180 a+148=180 a=180-148 a=32 So the angles are: first: 32 second: 78 third: 70